OpenStudy (anonymous):

Prove the law of sines using the cross product. Here is what I have done so far... A B C A+B+C=0 Reply Using Drawing From there I am stuck

OpenStudy (anonymous):



OpenStudy (jamesj):

Well, C x (A + B) = C x -C = 0 Thus C x A = - C x B C x A = B x C ...you take it from here

OpenStudy (anonymous):

From there you can express it as \[ACsin \theta = BCsin \phi \] C cancels and you are left with \[sin \theta /b = sin \phi /a\] Phi being the angle between B and C and theta being the angle between A and C. Yes?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

Thank you!